OpenStudy (anonymous):

Solve the equation. 1.3x+2.8=0.5x-4.48

OpenStudy (anonymous):

The first thing you should do is subtract 0.5x from both sides.

OpenStudy (anonymous):

That's so you can get rid of the 0.5x on the right side of the equation. 1.3x minus 0.5x is 0.7x.

OpenStudy (anonymous):

I got 0.8

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Allow me to check.

OpenStudy (anonymous):

\[1.3X + 2.8 = .5X - 4.48\]\[.8x + 2.8 = - 4.48\]\[.8x =-7.28\]\[x= -9.1\]

OpenStudy (anonymous):

You are correct, sorry. Then it should appear as this: 0.8x + 2.8 = -4.48 Add 4.48 to both sides. 0.8x = 7.28 Divide 0.8x with both sides to get rid of the 0.8x. You should get 9.1.

OpenStudy (anonymous):

How did you get -9.1?

OpenStudy (anonymous):

\[1.3(-9.10) + 2.8 = .5(-9.1) - 4.48\]

OpenStudy (anonymous):

i divided .8x with 7.28

OpenStudy (texaschic101):

it might make it a little easier if you start off by multiplying the entire equation by 100. This will get rid of the decimals. Just a suggestion

OpenStudy (texaschic101):

1.3x + 2.8 = 0.5x - 4.48 --- multiply everything by 100 130x + 280 = 50x - 448 -- subtract 50x from both sides 130x - 50x + 280 = -448 -- subtract 280 from both sides 130x - 50x = -448 - 280 -- combine like terms 80x = - 728 -- divide by 80 x = -728/80 x = -9.1

OpenStudy (anonymous):

I got it. I like the way the problem is set up. Thank You.